OpenStudy (danmac0710):

How can I integrate please: x/sqrt(1-x^2)

OpenStudy (amistre64):

consider a substitution? u = 1-x^2 or a trig sub? x = sin

OpenStudy (amistre64):

if x = sin, dx = cos dt sin cos ------------ dt sqrt(1-sin^2) sin cos ------------ dt cos sin dt

OpenStudy (danmac0710):

thanks amistre. You're wonderful!

OpenStudy (danmac0710):

How do you recognise when to use a trig sub? Do I just need to do a lot more practice?

OpenStudy (amistre64):

practice helps alot, the usub would have worked too u = 1-x^2 du = -2x dx -du/2 = x dx integrate -1/2sqrt(u) du = -sqrt(u) is the same resutls